OpenStudy (anonymous):

Help? I'm confused! How do I solve? |k+3|=6 |k-3|=6

OpenStudy (anonymous):

For solving absolute values, recall that: \[|a| = b \implies a = b \text{ or } a = -b\]

OpenStudy (saifoo.khan):

|k + 3| this MOD means that a variable will have 2 suspected values, it can be -ve or a +ve.

OpenStudy (anonymous):

So you just have to solve those two equations, and the union of the two solutions will be the solution set to the original

OpenStudy (anonymous):

k+3 = 6 or k+3 = -6 (using your first equation as an example)

OpenStudy (anonymous):

But how do i do it? do i subtract 3 from the first question?

OpenStudy (anonymous):

can you show me step by step for question number 1?

OpenStudy (anonymous):

No. You convert it to the two equations I have above, then solve each one separately. The solutions to the original equation will be the union of these two new equations.

OpenStudy (anonymous):

\begin{array}{ccc}k+3 = 6 & \text{ or } & k + 3 = -6\end{array} \[\implies \begin{array}{ccc}k= 3 & \text{ or } & k = -9\end{array}\]

OpenStudy (anonymous):

OH! I understand now! Thank You both!

OpenStudy (anonymous):

Either one of those two solutions will satisfy your original: |3+3| = 6 |-9 + 3| = 6